# From Champagne to Confined Polymer: Natural and Artificial Bubble Nucleation

### v2

@@ Abstract
In this study, we present a novel experimental work on bubble nucleation and growth using a model system comprised by viscoelastic polyvinyl butyral (PVB) confined in a Hele-Shaw cell geometry that is decompressed at elevated temperatures. The driver is connected to the temperature-induced shift in chemical equilibrium experienced simultaneously by two gases present in the bulk. The latter becomes at the same time oversaturated with water vapor and slightly undersaturated in air. Our bubbles grow with various shapes and sizes depending on the initial morphology of the nucleus or the presence of neighboring bubbles. The likelihood of nucleation is related to the amount of water dissolved in the bulk and the imposed temperature. Counter-intuitively, the number of nuclei whence a bubble can grow is inversely correlated with said temperature. In an analogy with Champagne, we show that nucleation can either be natural, at trapped fibers or dust particles, or artificial, at crenels we purposefully made in the glass surface. Our results indicate that the growth rate of bubbles can be impacted by the nucleation mechanism.
@@ I Introduction
Bubbles are magnificent objects that, for better or worse, are at the center of many important applications. For example, they are extremely valuable as ultrasound contrast agents in diagnostic imaging, show potential for drug delivery [1, 2], are used to probe the local rheology of soft matter materials [3], and may even dazzle the senses as in Champagne tasting [4, 5]. They can, at the same time, be extremely detrimental for they are at the root of injuries in decompression sickness [6, 7], the delamination of thin films [8], and can induce early failure—or themselves constitute failure—as seen in solar panels or architectural safety glass [9, 10]. In these examples, bubble nucleation and growth occur in a variety of materials - binary liquids, biomaterials, gels, polymers - so it is vital to understand the physical processes at play in materials more complex than a simple liquid if we are to maximize their positive potential while minimizing their harmful effects.
@@
Post decompression, we observe that non-coalescing bubbles may form with a morphology that is linked to the size and/or shape of the nuclei or the close presence of neighboring bubbles. Bubbles that stem from small nuclei grow with a circular shape provided they are well-separated from other bubbles. Bubbles that originate from large nuclei grow with a certain degree of anisotropy that is reminiscent of the original contour of the nuclei. Regardless of nuclei size/shape, bubbles may develop anisotropy when growing near other bubbles. Meanwhile, the likelihood of nucleation/growth is linked to the amount of water initially dissolved in the bulk or the imposed temperature. With respect to the latter, higher temperatures result in seemingly fewer nuclei whence bubbles may grow.
Overall, our results show that bubble nucleation can be subdivided into natural and artificial mechanisms in analogous fashion to Champagne’s. In the natural case, bubbles nucleate in trapped matter at the glass/polymer interface such as speckles or fibers. In the artificial case, bubbles grow from imperfections we on purpose made on the glass. While in both nucleation mechanisms bubbles exhibit the hallmark of diffusive growth, i.e., $R\sim(kt)^{1/2}$, where $R$ is the bubble radius, $k$ is the growth rate coefficient, and $t$ the time, we found that $k$ is affected by the nucleation mechanism itself. In the natural case, $k$ is globally the same for all bubbles which reflects in a way the total oversaturation of the polymer. In the artificial case, however, the growth rate is strongly influenced locally by the size of the crenel nucleus.
Setting aside our experimental conditions, we believe our results hold general truths for bubble nucleation and growth in composite, layered assemblies wherein gases or foreign matter become easily trapped at interfaces or in substrate defects. Two prime examples include laminated safety glass and photovoltaic modules, both staples of modern life wherein the PVB polymer is sandwiched in between two glass slides.
@@ II.2 Sample preparation
Conditioning. The polymer is stored under a controlled atmosphere to fix the amount of dissolved gases in the bulk, in particular air and water vapor. To this end, we stored the PVB polymer under humidity-regulated air at constant humidity activity $a$, either in a climatic chamber or in a desiccant box with silica gel for at least 48hrs. Three activity levels were used leading to the designations of wet ($a=0.5$), moist ($a=0.25$), and dry PVB ($a=0.05$) respectively. For simplicity, we treat air as nitrogen (N2, 78.1% in air) and disregard other gases (e.g., Ar, CO2) [^1] While oxygen can be found in relevant quantities in atmospheric air, its presence is of minimal relevance in our experiments. First, atmospheric air contains 3.71 times more N2 than O2—a dominating proportion when also considering that previous solubility studies have found that the PVB bulk absorbs only 1.5 times more oxygen than nitrogen [18]. Second, like N2, O2 is well above its critical point in our experiments, and thus, both gases can be treated as incondensable with an identical heat of solution sign [19, 20, 21]. When taken together, these two reasons imply that a bubble growing/dissolving in PVB with a mixture of N2/O2 will largely be driven by N2. Finally, the remaining gases in atmospheric air are present in negligible amounts and are likewise absorbed in very small quantities by the PVB polymer..
@@
Figure 1: Microscope images of polyvinyl butyral film when confined between two glass slides. A Original PVB surface roughness at room conditions. B Interfacial bubbles at the (top) glass/polymer interface after calendering. The shadowy regions are the unfocused bubbles from the bottom polymer/glass interface. B Experimental autoclave. C Sample schematics showing the imaging area (IA) whence the bubble measurements are taken.
@@ II.3 Bubble experiments
To study bubble nucleation and growth at high temperatures, we developed a transparent autoclave whose main body was constituted by a hollow cylinder made from stainless steel and a set of viewing windows (fig. 1C). A pair of heating ropes (Omegalux, 400 watts/unit, not shown) were wound around the periphery of the cylinder to supply heat, while pressure was provided via a Fluigent pressure generator (2 bar) connected to a dedicated port in the autoclave. Images were taken at the center of the glass sample in a target imaging area (IA) using an overhead high-resolution camera (UI-3240ML) with a 75mm (or 105mm) optical objective (res.$\sim 10-18.6\mu$m px-1), while light was shone from below the autoclave using an LED source (Schott KL 2500, not shown). During the experiments, the temperature was regulated using a PID controller and a thin film PT-100 sensor secured midway the glass sample at 15 mm away from the right edge (fig. 1D).
The experimental protocol included a heating ramp to soften the polymer under a hydrostatic load. The temperature was then held constant and the pressure rapidly decreased. In doing so, we increased the likelihood of bubble nucleation and growth by shifting the chemical equilibrium of the gases dissolved in the polymer bulk. Said shift is primarily linked to the variation in temperature from room conditions $T_{i}=25$°C to the temperature of decompression $T_{dec}$.
@@ III Results & Discussion
A bottle of Champagne contains six times its volume in terms of CO2. This is one of the secrets behind its famous effervescence: the liquid is oversaturated [4]. When the bottle is uncorked, the liquid wants to lose this excess gas to achieve chemical equilibrium with the surrounding air. Part of the CO2 in the liquid escapes via the free interface and the remaining gas is available to form bubbles. Our situation is somewhat analogous. During conditioning, we set the initial amount of gas in the polymer bulk. By heating and compressing/decompressing, we shift the chemical equilibrium of the gases in the bulk, and thus mimic the effect induced by uncorking in Champagne. In principle, bubbles should then grow (or not) depending on whether the PVB polymer wants to lose gas or absorb it at the temperature of decompression $T_{dec}$.
@@
Our post-decompression bubbles are inherently multi-component given the prevalence of air and water in the model system. Each gas has their respective oversaturation driver $f_{a}$, $f_{w}$. Contrary to Champagne (or other decompression bubbles in the literature [22, 23, 24, 25]), $f$ is in our case largely independent of the pressure drop (aside from possible entropic effects which we disregard) because our target area in the PVB polymer is closed off from the surrounding atmosphere given that (i) the glass is impermeable to our gases, and (ii) our samples are relatively large for the timescale set by diffusion. This also means that the PVB polymer absorbs an excess quantity of gas only in accordance with that which was initially present in the interfacial bubbles themselves. The latter are comprised mostly by air because of the conditioning atmosphere that was used, i.e., ${\color[rgb]{1,0,0}x_{a}}={\color[rgb]{1,0,0}n_{a}}/n\sim{\color[rgb]{1,0,0}p_{a}}/P_{o}\gtrsim 0.99$, where $x_{a}$, $n_{a}$, $p_{a}$ are the mol fraction, number of moles, and partial air pressure in the conditioning atmosphere, and $P_{o}$ is the reference atmospheric pressure (101.325 kPa).
We suppose the instantaneous driver $f$, before decompression, to be primarily given by the solubility ($H$) preferences of the gases in the PVB polymer. This is directly linked to their heat of solution $\Delta H_{s}$ and the shift in temperature. For water, $\Delta H_{s,w}$$<0$ [17], whereas for air $\Delta H_{s,a}$$>0$. See Appendix A for a gravimetric study of our own where we show that $N_{2}$ (the proxy for air) behaves remarkably close as in simpler, rubbery homopolymers such as polyisoprene in terms of sorption and diffusion.
@@
For the air driver, we must first compute the total amount of air in the system to then determine the effective solubility constant, and thus, the likelihood of air bubbles. We then have,
${\color[rgb]{1,0,0}c_{a}^{*}}={\color[rgb]{1,0,0}H_{a}}(T_{i}){\color[rgb]{1,0,0}p_{i,a}}+A\frac{e}{h}{\color[rgb]{1,0,0}\rho_{a}}(T_{i})$ (1)
wherein the left-hand side is the amount of air initially contained in the PVB bulk from natural saturation conditions and the right-hand side is the initial volume of air present in the interfacial bubbles per cubic volume of PVB, with $H_{a}$ being the solubility at the initial temperature $T_{i}$, $A$ is the initial fraction area of interfacial bubbles, typically $\approx 0.32\pm 0.02$SD in our experiments, $e$ is the characteristic polymer roughness, $h$ is the half-thickness of the PVB layer and $\rho_{a}$$(T_{i})$ is the dry air density included for dimensional homogeneity.
@@
Role of gas concentration. Figs. 2A,B compare the bubbling behaviour between dry and moist PVB. when using dry PVB, no bubbles nucleated after decompression in the time span of the experiment (fig. 2B). This meant that the air initially present in the anisotropic interfacial bubbles, and which was forcibly dissolved via the hydrostatic load, remained locked in solution inside the PVB bulk ($\Delta H_{s,a}$$>0$, $f_{a}$$\leq 1$). Contrastingly, when using moist PVB, bubbles nucleated and grew (fig. 2B), which is consistent with water vapour escaping from the PVB bulk with temperature ($\Delta H_{s,w}$$<0$, $f_{w}$$>1$). Thus, post decompression, observing either bubble growth (or not) is linked to the initial amount of water in the polymer bulk.
@@
Temperature effect. In figs. 2C-E, we surveyed the effect of temperature on bubble nucleation and growth at 100, 120, and 140°C. As seen therein, the number of bubbles post decompression and their size was dependent on $T_{dec}$. In this regard, fewer bubbles, albeit larger in size, were observed for 140°C (fig. 2D) compared to 120°C (fig. 2C).
@@
By continuing the latter experiment and increasing $T_{dec}$ from 100 to 120°C, we observed an explosive appearance of bubbles (fig. 2F). This suggests that the sample in fig. 2E had a large number of nuclei that only became active once we raised the temperature. This leads to a twofold conclusion: (i) the number of nucleation sites post decompression is apparently inversely proportional to $T_{dec}$ (as set by the number of bubbles in figs. 2C,D,F) and (ii) seemingly inactive sites can be activated by increasing the temperature. Most likely, the sites which we activated in fig. 2C by increasing the temperature to 120°C might have also become active at 100°C had we waited long enough. How long, however, is unknown.
While bubbles nucleated at seemingly different times, their growth history collapsed onto a single curve, as portrayed simultaneously in fig. 2E for 120,140°C, when arbitrary selecting $t_{o}=0$ s for $R_{o}\geq 50\mu$m. This selection of $t_{o}$ plays no role on the observed kinetics and any other bubble radius can be selected without affecting the results. We found that the growth curves of the bubbles at 120°C, 140°C are well described by a simple diffusion model [26] (dashed line in fig. 2G):
$R^{2}=kt+R_{o}^{2}$ (3)
@@
For the results provided in fig. 2G, the variation of the growth rate with temperature $k(140^{\circ}C)$/$k(120^{\circ}C)\sim 3.5$ is in the order of magnitude of the variation of $T_{dec}$ and the transport properties for water, which is respectively $\sim T_{140}/T_{120}\times D_{w}(140^{\circ}C)/D_{w}(120^{\circ}C)\times H_{w}(120^{\circ}C)/H_{w}(140^{\circ}C)\sim 3$ (see Tab. 4 in Appendix A for the transport constants). This confirms that our bubbles are globally driven by water vapour.
@@
Figure 3: A Artificial bubble nucleation and growth in moist PVB post decompression in four crenels of varying volume ($V_{crenel}$) etched on the glass surface ($T_{dec}\sim$144°C, $\Delta P=0.6$ bar). B $\circ$ Projected bubble radius $R$ versus time. The dashed line is a diffusion model $R\sim t^{1/2}$. The $R$ vs $t$ curves are color matched to the crenel volume in A. $\circ$ Bubble growth from natural nucleation. The inset plots the growth rate $k$ versus crenel volume. Symbols and solid line are color matched to their respective bubbles. The (teal) horizontal baseline is the growth rate obtained from a naturally nucleated bubble. The dashed black line is a linear regression.
@@
Bubble morphology. In Champagne, bubbles detach from their nucleation site and rapidly become spherical and ascend by buoyancy. The bubbles maintain sphericity because the fluid is newtonian. In non-newtonian fluids or brittle materials, by contrast, it has been shown that bubbles can take on complex shapes when the rate of deformation is larger than the polymer relaxation time [30, 31]. Similarly, in a hydrogel, neighboring bubbles develop curved segments towards one another, whenever the distance separating them is roughly equivalent to their respective radii [23]. In our case, the initial bubble morphology is linked to the shape of the nuclei before decompression or the presence of neighboring bubbles.
@@
Figure 4: Morphology of bubbles post decompression ($T_{dec}\sim 140$°C, $\Delta P=0.6$ bar). A-B Undissolved anisotropic bubbles (moist PVB) serve as nuclei for bubble growth. The latter grow with a contour resembling the original shape of the nucleus before approaching circularity as time passes. C Bubbles with an initial circular projection develop anisotropy when sufficiently close to neighboring bubbles (the sequence has been shifted by one minute to coincide with the scale from A-B). D A collection of partially dissolved anisotropic bubbles lead to irregular bubble morphologies (PVB conditioning a=0.5). The anisotropy is not linked to the high water content (not shown) but rather to the presence of closely-packed anisotropic bubbles. E-F Examples of non-coalescing bubbles. In E, bubbles grew close to each other at the same polymer/glass interface and deformend as they came in close proximity. In F, three bubbles grew at different interfaces. The small bubbles grew at the bottom interface, while the large one is attached to the top interface.

### v3

@@ Abstract
In this study, we present an experimental work on bubble nucleation and growth using a model system comprised of viscoelastic polyvinyl butyral confined in a Hele-Shaw cell geometry that is decompressed at elevated temperatures. The appearance and growth of bubbles are connected to the temperature-induced shift in chemical equilibrium experienced simultaneously by two gases present in the bulk. The latter becomes simultaneously oversaturated with water vapor and slightly undersaturated in air. Our bubbles grow with various shapes and sizes depending on the initial morphology of the nucleus or the presence of neighboring bubbles. For large nuclei, bubbles grow anisotropically because of contact line pinning. The likelihood of nucleation is related to the amount of water dissolved in the bulk and the imposed temperature. Counter-intuitively, the number of nuclei whence a bubble can grow is inversely correlated with said temperature. In an analogy with champagne, we show that nucleation can either be natural, at trapped fibers or dust particles, or artificial, at crenels we purposefully made in the glass surface. Our results indicate that the growth rate of bubbles can be impacted by the nucleation mechanism.
@@ I Introduction
Bubbles are magnificent objects that, for better or worse, are at the center of many important applications. For example, they are extremely valuable as ultrasound contrast agents in diagnostic imaging, show potential for drug delivery [1, 2], are used to probe the local rheology of soft matter materials [3], and may even dazzle the senses as in champagne tasting [4, 5]. They can, at the same time, be extremely detrimental for they are at the root of injuries in decompression sickness [6, 7], the delamination of thin films [8], and can induce early failure—or themselves constitute failure—as seen in solar panels or architectural safety glass [9, 10]. In these examples, bubble nucleation and growth occur in a variety of materials - binary liquids, biomaterials, gels, polymers. In such non-newtonian fluids, bubbles exhibit particular behaviors such as interaction between neighboring bubbles, which may deform toward one another, as if attracting each other [23], or may develop irregular shapes when the local deformation rate exceeds the relaxation time of the polymer [30, 31]. It is therefore important to understand the physical processes at play in materials more complex than a simple liquid if we are to maximize their positive potential while minimizing their harmful effects.
@@
Post decompression, we observe that non-coalescing bubbles may form with a morphology that is linked to the size and/or shape of the nuclei or the close presence of neighboring bubbles. Bubbles that stem from small nuclei grow with a circular shape provided they are well-separated from other bubbles. Bubbles that originate from large nuclei grow with a certain degree of anisotropy that is reminiscent of the original contour of the nuclei because the initial contact line of the bubble remains pinned. Regardless of nuclei size/shape, bubbles may develop anisotropy when growing near other bubbles. Meanwhile, the likelihood of nucleation/growth is linked to the amount of water initially dissolved in the bulk or the imposed temperature. With respect to the latter, higher temperatures —wherein the polymer becomes softer—result paradoxically in seemingly fewer nuclei whence bubbles may grow.
Overall, our results show that bubble nucleation can be subdivided into natural and artificial mechanisms in analogous fashion to champagne’s. In the natural case, bubbles nucleate in trapped matter at the glass/polymer interface such as speckles or fibers. In the artificial case, bubbles grow from imperfections we purposely made on the glass. While in both nucleation mechanisms bubbles exhibit the hallmark of diffusive growth, i.e., $R\simeq(kt)^{1/2}$[26], where $R$ is the bubble radius, $k$ is the growth rate coefficient, and $t$ the time, we found that $k$ is affected by the nucleation mechanism itself. In the natural case, $k$ is globally the same for all bubbles, which reflects in a way the total oversaturation of the polymer. In the artificial case, however, the growth rate is strongly influenced locally by the size of the crenel nucleus.
Setting aside our experimental conditions, we believe our results hold general truths for bubble nucleation and growth in composite, layered assemblies wherein gases or foreign matter become easily trapped at interfaces or in substrate defects. Two prime examples include laminated safety glass and photovoltaic modules, both staples of modern life wherein the PVB polymer is sandwiched between two glass slides and where large variations in temperature are necessary to bond the polymer to the confining glass.
@@ II.2 Sample preparation
Conditioning. The polymer is stored under a controlled atmosphere to fix the amount of dissolved gases in the bulk, in particular air and water vapor. To this end, we stored the PVB polymer under humidity-regulated air at constant activity $\varphi=p_{w}/P_{Sat}(T)$, where $p_{w}$ is the vapour pressure in the atmosphere and $P_{Sat}$ is the saturation pressure at the temperature $T$. This was performed in either a climatic chamber or in a desiccant box with silica gel for at least 48hrs. Three activity levels were used leading to the designations of wet ($\varphi=0.5$), moist ($\varphi=0.25$), and dry PVB ($\varphi=0.05$) respectively. For simplicity, we treat air as nitrogen (N2, 78.1% in air) and disregard other gases (e.g., Ar, CO2) [^1] We chose to work experimentally with atmospheric air because this gas is prevalent during the manufacturing of architectural glass and photovoltaic panels. While oxygen can be found in relevant quantities in the aforementioned gas, its presence is of minimal relevance in our experiments. First, atmospheric air contains 3.71 times more N2 than O2—a dominating proportion when also considering that previous solubility studies have found that the PVB bulk absorbs only 1.5 times more oxygen than nitrogen [18]. Second, like N2, O2 is well above its critical point in our experiments, and thus, both gases can be treated as incondensable with an identical heat of solution sign [19, 20, 21]. When taken together, these two reasons imply that a bubble growing/dissolving in PVB with a mixture of N2/O2 will largely be driven by N2. Finally, the remaining gases in atmospheric air are present in negligible amounts and are likewise absorbed in very small quantities by the PVB polymer..
@@
Figure 1: Microscope images of polyvinyl butyral film when confined between two glass slides. A Original PVB surface roughness at room conditions. B Interfacial bubbles at the (top) glass/polymer interface after calendering. The shadowy regions are the unfocused bubbles from the bottom polymer/glass interface. C Experimental autoclave. D Sample schematics showing the imaging area (IA) whence the bubble measurements are taken.
@@ II.3 Bubble experiments
[…]
The experimental protocol included the application of a heating ramp under a hydrostatic load, followed by rapid decompression of the autoclave chamber under isothermal conditions. The hydrostatic load was applied to dissolve the interfacial bubbles initially present in the sample (see fig. 1B) inside the PVB polymer bulk. Meanwhile, the variation in temperature, from room conditions $T_{i}=25$°C to the temperature of decompression $T_{dec}$, triggered the nucleation and growth of bubbles post decompression by shifting the chemical equilibrium of the gases (air and water) dissolved in the polymer bulk.
@@ III Results & Discussion
A bottle of champagne contains six times its volume in terms of CO2. This is one of the secrets behind its famous effervescence: the liquid is oversaturated [4]. When the bottle is uncorked, the liquid releases this excess gas to achieve chemical equilibrium with the surrounding air. Part of the CO2 in the liquid escapes via the free interface and the remaining gas is available to form bubbles. Our situation is somewhat analogous. During conditioning, we set the initial amount of gas in the polymer bulk. By heating and compressing/decompressing, we shift the chemical equilibrium of the gases in the bulk, and thus mimic the effect induced by uncorking in champagne. In principle, bubbles should then grow (or not) depending on whether the PVB polymer releases or absorbs gases at the decompression temperature $T_{dec}$.
@@
Relative gas saturation. Our post-decompression bubbles are inherently multi-component given the prevalence of air and water in the model system. Each gas has their respective relative saturation $f_{a}$, $f_{w}$ that is set independently by $f=1+\zeta$, where $\zeta=\Delta c/c_{0}$, and $c_{0}$ is the concentration at saturation conditions (room conditions in our case) and $\Delta c$ is the concentration difference relative to $c_{0}$ at the decompression temperature $T_{dec}$. Broadly speaking, $f>1$ means gas oversaturation and is linked to bubble growth, $f=1$ is the reference saturation state wherein gases are in equilibrium and bubbles are in principle static (whenever surface tension or other rheological effects are negligible), and $f<1$ implies gas undersaturation and concomitantly bubble shrinkage.
Contrary to other decompression bubbles in the literature [22, 23, 24, 25], $f$ is in our case largely independent of the pressure drop (aside from possible entropic effects which we disregard) because our target area in the PVB polymer is closed off from the surrounding atmosphere given that (i) the confining glass is impermeable to our gases, and (ii) our samples are relatively large for the timescale set by diffusion. This also means that during our experiments the PVB polymer absorbs an excess quantity of gas only in accordance with that which was initially present in the interfacial bubbles themselves. The latter are comprised mostly of air because of the conditioning atmosphere that was used, i.e., ${\color[rgb]{1,0,0}x_{a}}={\color[rgb]{1,0,0}n_{a}}/n\sim{\color[rgb]{1,0,0}p_{a}}/P_{o}\gtrsim 0.99$, where $x_{a}$, $n_{a}$, $p_{a}$ are the mol fraction, number of moles, and partial air pressure in the conditioning atmosphere, and $P_{o}$ is the reference atmospheric pressure (101.325 kPa).
We suppose the instantaneous relative saturation $f$, before decompression, to be primarily influenced by how the Henry solubility constant ($H$, [$kg\ m^{-3}\ Pa^{-1}$]) of the gases evolves during our experiments. This is directly linked to their heat of solution $\Delta H_{s}$ in the PVB polymer and the shift in temperature from room conditions to $T_{dec}$. For water, $\Delta H_{s,w}$<0<0< 0 [17], whereas for air $\Delta H_{s,a}$>0>0> 0. See the Appendix for a gravimetric study of our own where we show that $N_{2}$ (the proxy for air) behaves remarkably close as in simpler, rubbery homopolymers such as polyisoprene in terms of sorption and diffusion.
@@
For the relative air saturation, we must first compute the total concentration of air in the system to then determine the effective solubility constant, and thus, the likelihood of air bubbles. Accordingly, we define the total air concentration ${\color[rgb]{1,0,0}c_{a}^{*}}$ [$kg\ m^{-3}$],
${\color[rgb]{1,0,0}c_{a}^{*}}={\color[rgb]{1,0,0}c_{d}}+{\color[rgb]{1,0,0}c_{i}},$ (1)
where ${\color[rgb]{1,0,0}c_{d}}$ is the concentration of air initially dissolved in the PVB bulk from natural saturation conditions and ${\color[rgb]{1,0,0}c_{i}}$ is the air concentration set by the initial volume of air present in the interfacial bubbles per unit volume of PVB.
${\color[rgb]{1,0,0}c_{d}}$ is given by
${\color[rgb]{1,0,0}c_{d}}={\color[rgb]{1,0,0}H_{a}}(T_{i}){\color[rgb]{1,0,0}p_{i,a}},$ (2)
with $H_{a}$ being the solubility constant at the initial temperature $T_{i}$ while ${\color[rgb]{1,0,0}p_{i,a}}$ is the initial air partial pressure (conditioning pressure).
Meanwhile, ${\color[rgb]{1,0,0}c_{i}}$ is determined as
${\color[rgb]{1,0,0}c_{i}}=A\frac{e}{h}{\color[rgb]{1,0,0}\rho_{a}}(T_{i}),$ (3)
wherein $A$ is the initial fraction area of interfacial bubbles, typically 0.32±0.02$\approx$ 0.32$\pm$ 0.02≈ 0.32 ± 0.02SD in our experiments, $e$ is the characteristic polymer roughness, $h$ is the half-thickness of the PVB layer and $\rho_{a}$$(T_{i})$ is the dry air density included for dimensional homogeneity with ${\color[rgb]{1,0,0}c_{d}}$.
@@
Role of gas concentration. Figs. 2A,B compare the bubbling behaviour between dry and moist PVB. When using dry PVB, no bubbles nucleated after decompression in the time span of the experiment (fig. 2A). This meant that the air initially present in the anisotropic interfacial bubbles, and which was forcibly dissolved via the hydrostatic load, remained locked in solution inside the PVB bulk ($\Delta H_{s,a}$>0>0> 0, $f_{a}$1$\leq$ 1≤ 1). Contrastingly, when using moist PVB, bubbles nucleated and grew (fig. 2B), which is consistent with water vapour escaping from the PVB bulk with temperature ($\Delta H_{s,w}$<0<0< 0, $f_{w}$>1>1> 1). Thus, post decompression, observing either bubble growth (or not) is linked to the initial amount of water in the polymer bulk.
@@
Temperature effect. In figs. 2C-E, we surveyed the effect of temperature on bubble nucleation and growth at 100, 120, and 140°C. As seen therein, the number of bubbles post decompression and their size was dependent on $T_{dec}$. In this regard, several bubbles, albeit smaller in size but in larger numbers, were observed at 120°C (fig. 2C) compared to 140°C (fig. 2D).
@@
By continuing the latter experiment and increasing $T_{dec}$ from 100 to 120°C, we observed an explosive appearance of bubbles (fig. 2F). This suggests that the sample in fig. 2E had a large number of nuclei that only became active once we raised the temperature [^2] We visualized our samples under a microscope to qualitatively compare the number of nuclei therein as a function of $T_{dec}$. To this end, we sampled the number of gas inclusions at the polymer/glass interfaces but disregarded the bubbles already detected by our autoclave system. Indeed, the number of said inclusions was inversely correlated with $T_{dec}$. However, ascertaining the quantitative relationship between the two was beyond the scope of the present work.. This leads to a twofold conclusion: (i) the number of nucleation sites post decompression is apparently negatively correlated with $T_{dec}$ (as set by the number of bubbles in figs. 2C,D,F) and (ii) seemingly inactive sites can be activated by increasing the temperature. Most likely, the sites which we activated in fig. 2E by increasing the temperature to 120°C might have also become active at 100°C had we waited long enough. How long, however, is unknown.
We speculate there are two effects at play for the paradoxical relationship between $T_{dec}$ and the number of bubbles/nuclei. First, between 100-140°C, the rheology of the polymer varies significantly. For comparison, we estimate the shear modulus at 100°C ($t=1$s) to be $G\sim 10^{5}$ Pa, which then decreases by an order of magnitude to $G\sim 10^{4}$ Pa ($t=10^{2}$s). Contrastingly, at 140°C, the PVB polymer is between 3-15 times correspondingly softer [15]. We thus hypothesize that the high stiffness at 100°C probably restrained the nuclei from growing into visible bubbles, despite the relatively high water saturation ($\color[rgb]{0,0,1}{f_{w}}$ 40$\sim$ 40∼ 40). Second, the higher the temperature, the more easily air can dissolve in the PVB bulk, thus diminishing the number of nuclei whence a readily visible bubble may eventually grow. More work is needed to conclusively verify the mechanism behind the number of nuclei that grow into visible bubbles.
Diffusion model. While bubbles nucleated at seemingly different times, their growth history collapsed onto a single curve, as portrayed simultaneously in fig. 2G for 120,140°C, when arbitrarily defining $t_{o}=0$ s for $R_{o}\geq 50\mu$m. This selection of $t_{o}$ plays no role on the observed kinetics and any other bubble radius can be selected without affecting the results. We found that the growth curves of the bubbles at 120°C, 140°C are well described by a simple diffusion model [26] (dashed line in fig. 2G):
$R^{2}=kt+R_{o}^{2},$ (5)
@@
For the results provided in fig. 2G, the variation of the growth rate with temperature $k(140^{\circ}C)$/$k(120^{\circ}C)\sim 3.5$ is in the order of magnitude of the variation of $T_{dec}$ and the transport properties for water, which is respectively $\sim T_{140}/T_{120}\times D_{w}(140^{\circ}C)/D_{w}(120^{\circ}C)\times H_{w}(120^{\circ}C)/H_{w}(140^{\circ}C)\sim 3$ (see tables  1-4 in the Appendix for the transport constants). This confirms that our bubbles are globally driven by water vapour.
@@
Figure 3: A Artificial bubble nucleation and growth in moist PVB post decompression in four crenels of varying volume ($V_{crenel}$) etched on the glass surface ($T_{dec}\sim$144°C, $\Delta P=0.6$ bar). B $\circ$ Blue, red, and purple circles indicate projected bubble radius $R$ versus time. The dashed line is a diffusion model $R\sim t^{1/2}$. The $R$ vs $t$ curves are color matched to the crenel volume in A. $\circ$ Teal circles indicate bubble growth from natural nucleation. The inset plots the growth rate $k$ versus crenel volume. Symbols and solid line are color matched to their respective bubbles. The (teal) horizontal baseline is the growth rate obtained from a naturally nucleated bubble. The dashed black line is a linear regression.
@@
Bubble morphology. In champagne, bubbles detach from their nucleation site and rapidly become spherical and ascend by buoyancy. The bubbles maintain sphericity because they are smaller than the capillary length and the fluid is newtonian. In non-newtonian fluids or brittle materials, by contrast, it has been shown that bubbles can take on complex shapes when the rate of deformation is larger than the polymer relaxation time [30, 31]. Similarly, in a hydrogel, neighboring bubbles develop curved segments towards one another, whenever the distance separating them is roughly equivalent to their respective radii [23]. In our case, the initial bubble morphology is linked to the shape of the nuclei before decompression or the presence of neighboring bubbles.
Figure 4: Morphology of bubbles post decompression ($T_{dec}\sim 140$°C, $\Delta P=0.6$ bar). A-B Undissolved anisotropic bubbles (moist PVB) serve as nuclei for bubble growth. The latter grow with a contour resembling the original shape of the nucleus before approaching circularity as time passes. C Bubbles with an initial circular projection develop anisotropy when sufficiently close to neighboring bubbles (the sequence has been shifted by one minute to coincide with the scale from A-B). D A collection of partially dissolved anisotropic bubbles lead to irregular bubble morphologies (PVB conditioning $\varphi=0.5$). The anisotropy is not linked to the high water content (not shown) but rather to the presence of closely-packed anisotropic bubbles. E-F Examples of non-coalescing bubbles. In E, bubbles grew close to each other at the same polymer/glass interface and deformend as they came in close proximity. In F, three bubbles grew at different interfaces. The small bubbles grew at the bottom interface, while the large one is attached to the top interface.